OpenStudy (anonymous):

1/8+3/4 a.1/2 b.1/3 c.1/4 d.2/3

OpenStudy (anonymous):

3/4=6/8, so 1/8+6/8=7/8, so none of those

OpenStudy (anonymous):

estimate

OpenStudy (anonymous):

the closest would be 2/3